OpenStudy (anonymous):

laplace transform: L{sin^3 (at)} so for I have this: L{ sin(at) * [1/2 ( 1 - cos(2at)]} 1/2 L{sin(at)} - 1/2 L{cos(2at)*sin(at)} 1/2 * (a^2)/(s^2 + a^2) - 1/2 L{cos(2at)*sin(at)} How do i get rid of the cos(2at)*sin(at) ?

OpenStudy (ybarrap):

try converting the trig functions to their complex equivalents (e^(ix))

OpenStudy (anonymous):

using Eulers formula does the job, posting the work in a sec.

OpenStudy (anonymous):

im trying to get a trignometric relation with sin^3(at) instead of doing sin(at)*sin^2(at) = sin(at)*(1/2 - 1/2 * cos(2at)

OpenStudy (anonymous):

wtf did you just make happen there? Huge thanks! I was trying to use this explanation http://answers.yahoo.com/question/index?qid=20080723124442AAcZCTX but i didnt get that either. Anyway, thanks for the relation

OpenStudy (anonymous):

Here are variations of Euler's Formulas that may help with problems like these: \[e^{i\theta}=\cos\theta+i \sin\theta\] \[\cos\theta=\frac{e^{i\theta}+e^{-i\theta}}{2}\] \[\sin\theta=\frac{e^{i\theta}-e^{-i\theta}}{2i}\]

OpenStudy (anonymous):

oh now i get it. Thanks sir

OpenStudy (anonymous):

help pls....Laplace transform: L{cos^3 (at)}